OpenStudy (anonymous):

If √6 is the geometric mean between 4 and another number, then the number is A.1.5 B.9 C.24

OpenStudy (anonymous):

geometric mean of \(a,b\) is \(\sqrt{ab}\) you are told \(\sqrt{4\times b}=\sqrt{6}\) so \(b=\frac{6}{4}\)

OpenStudy (anonymous):

Okay thank you so much